0.2 g of a sample of H2O2 is


required for 10 ml of 1 N KMnO4 in a titration in the presence of H2SO4. Purity of H2O2



is (a) 25% (b) 85% (c) 65% (d) 95%


Grade:12

1 Answers

Vikas TU
10 years ago

In 0.2 g of a sample of H2O2,

Let ''x'' gm of pure H2O2 is present, then

Equivalents of H2O2 = Equivalents of KMnO4

moles of H202 X V.F of H202 = moles of KMnO4 X V.F of KMnO4

           ''''                                       = Molarity x volume x V.f of KMnO4

           ''''                                       = Normality x V.F of KMnO4   [N =M . V.f]

   thus,

  

(x/34) x 2 = 1 x 10/1000

x/17 = 1/100

x = 17/100

x = 0.17.

 

Thus Pure H202 in 0.2 gm sample is =

0.17/0.2 x 100

= 85 %

I hpe this correct!

Whenever titration prob. comes try to use equivalents of each ractants and products if it is!

Approve!